OpenStudy (anonymous):

During a sale, a dress decreased in price from $60 to $48. What was the percent decrease in the price of the dress?

OpenStudy (anonymous):

\[\frac{ 48 }{ 60 }\times 100\]

OpenStudy (bibby):

If you need a general formula. I believe it goes (older-newer)/older 60-48 = 12 12/60 = 0.2 = 20% or you can get the ratio 48/60 and multiply by 100